OpenStudy (theopenstudyowl):

Was wondering if someone could go over these homework answers with me before I turn this in.

OpenStudy (theopenstudyowl):

Which of the following represents the information you need in order to find the sample size necessary for a confidence interval for a single population proportion with a given margin of error and confidence level? z*, n, t*, m, p* t*, n, p* z*, m, p* z*, m, s

OpenStudy (theopenstudyowl):

I think it's C

OpenStudy (michele_laino):

sorry can you write the meaning of z, m, s, t, p?

OpenStudy (theopenstudyowl):

sure, one sec...

OpenStudy (theopenstudyowl):

Z or z refers to a standardized score, also known as a z score.

OpenStudy (michele_laino):

ok!

OpenStudy (theopenstudyowl):

its ok, I think I known the answe to this one, could I ask another one in the same thread??

OpenStudy (michele_laino):

ok!

OpenStudy (theopenstudyowl):

A company wants to find out what kinds of transportation its employees use to get to work. It conducts a survey of 537 employees, and 243 say they ride the bus. Construct a 95% confidence interval for the proportion of employees who ride the bus to and from work. (.411, .495) (.4521, .4539) (2.168, 2.252) (.3962, .5098) (.4316, .4744) I think it's C

OpenStudy (anonymous):

I'm sure it's b

OpenStudy (theopenstudyowl):

prove it

OpenStudy (michele_laino):

we have a probability 243/537 for going with bus, and a probability of 1-(243/537) for not going with bus

OpenStudy (michele_laino):

oops...1- (243/537)

OpenStudy (michele_laino):

I think that we have to use the binomial distribution

OpenStudy (michele_laino):

so the standard deviation is: \[\Large \sigma = \sqrt {Npq} = \sqrt {537\frac{{243}}{{537}}\left( {1 - \frac{{243}}{{537}}} \right)} = ...\]

OpenStudy (michele_laino):

whereas the mean is: \[\Large m = Np = 243\]

OpenStudy (michele_laino):

then we can write: \[\begin{gathered} m = Np = 243 \hfill \\ \hfill \\ \sigma = 11.534 \hfill \\ \end{gathered} \]

OpenStudy (theopenstudyowl):

ok

OpenStudy (michele_laino):

next, a confidence level of 95% mean a distance about 1.96* sigma

OpenStudy (theopenstudyowl):

oh ok

OpenStudy (theopenstudyowl):

I think my answer is right! Thanks

OpenStudy (michele_laino):

ok! Thanks!

OpenStudy (theopenstudyowl):

is it right?

OpenStudy (theopenstudyowl):

did u get the same result?

OpenStudy (michele_laino):

please I continue then

OpenStudy (theopenstudyowl):

ok

OpenStudy (theopenstudyowl):

sorry

OpenStudy (michele_laino):

referring to our sample of population, we have: mean value= 243/537=0.452 and 1.96*sigma= 1.96*(11.534/537)=0.04 so our interval, is: 0.452-0.04=0.412 left value 0.452+0.04= 0.492 right value, or \[\Large \left( {0.412,\quad 0.492} \right)\]

OpenStudy (theopenstudyowl):

huh, looks like the closest answer is A???

OpenStudy (michele_laino):

yes!

OpenStudy (theopenstudyowl):

wow! thanks! And thnxs for walking me through the steps

OpenStudy (michele_laino):

thanks! :)